OpenStudy (anonymous):

What is the logarithmic function modeled by the following table? x f(x) 8 3 16 4 32 5 f(x) = logx2 f(x) = log2x f(x) = 2 log10x f(x) = x log102

jhonyy9 (jhonyy9):

what is your opinion please ?

OpenStudy (anonymous):

I have no idea on how to do these :(

jhonyy9 (jhonyy9):

log2 2 = 1 do you know this ?

jhonyy9 (jhonyy9):

and 2log2 2 = log2 2^2 = 2 and this you know ?

OpenStudy (anonymous):

nope... to be honest i was never thought about this

jhonyy9 (jhonyy9):

ok. 8= 2^? how many will be the exponent of 2 for what you will get 8 ?

jhonyy9 (jhonyy9):

come en please this is easy

jhonyy9 (jhonyy9):

2*2*2 = 8 so than 2^? how many will be the exponent ?

OpenStudy (anonymous):

2^3=8

jhonyy9 (jhonyy9):

exactly - nice

OpenStudy (anonymous):

ok

jhonyy9 (jhonyy9):

so and 2^? = 16 how many will be the exponent of 2 in this case ?

OpenStudy (anonymous):

2^4=16

OpenStudy (anonymous):

2^5=32

jhonyy9 (jhonyy9):

exactly now you need checking there from these choices what is logarithm base 2 of x do you see it anywhere ?

OpenStudy (anonymous):

is it b

jhonyy9 (jhonyy9):

write me please here to be sure that i understand you correct

OpenStudy (anonymous):

f(x)=log2x

OpenStudy (anonymous):

it is a base 2

jhonyy9 (jhonyy9):

vera vera nive - great job - hope so much that you have understood it really true yes ?

jhonyy9 (jhonyy9):

yes because after log the first number always sign the base

OpenStudy (anonymous):

i do... thank so much... can u help with some other questions please?

jhonyy9 (jhonyy9):

yes but please write them on the open questions column and tag me please there - ok ?

jhonyy9 (jhonyy9):

was my pleasure good luck bye

OpenStudy (anonymous):

ok